RC Meeting, Wednesday, October 07 , 2021

From HallCWiki
Jump to: navigation, search


RC Meeting Information

RC Daily Meetings are at 10 am, remote https://bluejeans.com/861439920

Accelerator Status

  1. Smooth beam delivery.
  2. Helicity board will be switched to no delay reporting and remain at 120Hz.
  3. Hall A likely to be ready to receive beam tomorrow mid-day. Experts are on standby to execute a fast setup for all 4 halls.

Hall C Status

  1. Beam delivery during the BCM calibration went smoothly ([1])
  2. Smooth running for the rest of the evening and over night.
  3. At 1:30 last night the 15K supply temperature shot up to 16.5 (https://logbooks.jlab.org/entry/3919367). This resulted in no beam for ~20min. Experts are looking into it.
  4. Dave M pointed out that our ice ball over the target has been growing since Sunday (https://logbooks.jlab.org/entry/3919436).

Current Run Plan

  1. Continue stable data taking until configuration change (hopefully over the weekend)
    • Production data on LH2 target with 45 uA beam at kinematics (SHMS: 8.035 GeV/c, 5.505 deg | HMS: -1.026 GeV/c, 47.07 deg)

Current Issues

  1. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.

Special Requests

Useful Links